# RLCP: A Reinforcement Learning-based Copyright Protection Method for Text-to-Image Diffusion Model

### v2

@@ Abstract
The increasing sophistication of text-to-image generative models has led to complex challenges in defining and enforcing copyright infringement criteria and protection. Existing methods, such as watermarking and dataset deduplication, fail to provide comprehensive solutions due to the lack of standardized metrics and the inherent complexity of addressing copyright infringement in diffusion models. To deal with these challenges, we propose a Reinforcement Learning-based Copyright Protection(RLCP) method for Text-to-Image Diffusion Model, which minimizes the generation of copyright-infringing content while maintaining the quality of the model-generated dataset. Our approach begins with the introduction of a novel copyright metric grounded in copyright law and court precedents on infringement. We then utilize the Denoising Diffusion Policy Optimization (DDPO) framework to guide the model through a multi-step decision-making process, optimizing it using a reward function that incorporates our proposed copyright metric. Additionally, we employ KL divergence as a regularization term to mitigate some failure modes and stabilize RL fine-tuning. Experiments conducted on 3 mixed datasets of copyright and non-copyright images demonstrate that our approach significantly reduces copyright infringement risk while maintaining image quality.
@@ 1 Introduction
Recently, text-to-image diffusion models have garnered significant attention in research. These advanced methods (Balaji et al. 2023; Nichol et al. 2022; Rombach et al. 2022a; Saharia et al. 2022) have demonstrated exceptional capabilities in converting textual descriptions into highly accurate and visually coherent images. The advancements in these techniques have unlocked numerous possibilities for various downstream tasks, including image editing (Avrahami, Lischinski, and Fried 2022; Ho, Jain, and Abbeel 2020a; Kawar et al. 2023), image denoising (Ho, Jain, and Abbeel 2020a; Xie et al. 2023), and super-resolution (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020b).
While the progress in text-to-image generative models has profoundly impacted different industries, it also presents significant challenges for copyright protection. These models utilize extensive training data that may include copyrighted works, which they are sometimes capable of memorizing(Carlini et al. 2023). This ability can result in the production of images that closely resemble protected content (See in Figure 1), posing significant challenges to copyright protection(Elkin-Koren et al. 2023). Recent legal cases, such as those involving Stable Diffusion(Rombach et al. 2022b) and Midjourney(Mansour 2023), highlight concerns over the use of copyrighted data in AI training, where the models potentially infringe on the rights of numerous artists. These cases highlight a growing concern: Could the high-quality content synthesized by these generative AIs be excessively similar to copyrighted training data, potentially violating the rights of copyright holders?
Various methods have been proposed for source data copyright protection. One approach involves using unrecognizable examples(Gandikota et al. 2023; Zhang et al. 2023) that prevent models from learning key features of protected images either during inference or training stages. However, this method is highly dependent on the specific image and model, and it lacks general reliability. Watermarking(Dogoulis et al. 2023; Epstein et al. 2023) inserts specific, unnoticeable patterns into protected images to detect copyright infringement, but further research is needed to improve its robustness. Machine unlearning(Bourtoule et al. 2020; Ginart et al. 2019; Huang et al. 2021; Gao et al. 2023; Nguyen et al. 2022) removes contributions of copyright data, aligning with the right to be forgotten, while dataset deduplication(Somepalli et al. 2022) helps reduce the risk of training sample memorization.
@@
To tackle these challenges, we propose a Reinforcement Learning-based Copyright Protection method (RLCP) for text-to-image diffusion models to reduce the possibility of generating copyright-infringing content. Specifically, inspired by Courts111https://law.justia.com/cases/federal/appellate-courts/F2/562/1157/293262/ in the US which employs two-part test to determine copyright violation, which contains an extrinsic test examining objective similarity in specific expressive elements, and an intrinsic test assessing subjective similarity from the perspective of a reasonable audience, we first propose copyright metric that mirror these legal standards by combining semantic and perceptual similarity. Then, we propose a novel framework that combines reinforcement learning with copyright infringement metrics. We leverage the Denoising Diffusion Policy Optimization (DDPO) framework to guide the model through a multi-step decision-making process, optimizing it using a reward function that incorporates our proposed copyright metric. Additionally, we employ KL divergence as a regularization term to mitigate some failure modes and stabilize RL fine-tuning.
@@ 2 Related Work
We briefly review related work about text-to-image diffusion models, copyright protection and Reinforcement Learning from Human Feedback.
@@
Text-to-Image Diffusion Models: Recently, text-to-image diffusion models have garnered significant attention in research. These advanced methods (Balaji et al. 2023; Nichol et al. 2022; Rombach et al. 2022a; Saharia et al. 2022) have demonstrated exceptional capabilities in converting textual descriptions into visually coherent and realistic images with high accuracy. The advancements in these techniques have unlocked numerous possibilities for various downstream tasks, including image editing (Avrahami, Lischinski, and Fried 2022; Ho, Jain, and Abbeel 2020a; Kawar et al. 2023), image denoising (Ho, Jain, and Abbeel 2020a; Xie et al. 2023), and super-resolution (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020b).
@@
Copyright Protection: Several studies in the legal literature have examined copyright issues in machine learning and data mining, focusing primarily on potential infringements during the training phase.: (1)Watermarking(Dogoulis et al. 2023; Epstein et al. 2023), which inserts specific, unnoticeable patterns into protected images to detect copyright infringement, has been explored, but further research is needed to improve its robustness. (2)Concept Removal: To remove explicit artwork from large models, (Gandikota et al. 2023) presents a fine-tuning method for concept removal from diffusion models. Additionally, (Zhang et al. 2023) presents the ”Forget-Me-Not” method, which enables the targeted removal of specific objects and content from large models within 30 seconds while minimizing the impact on other content. (3) Dataset Deduplication: (Somepalli et al. 2022) explores whether diffusion models create unique artworks or directly replicate certain content from the training dataset during image generation. (4) Machine Unlearning: Numerous model unlearning methods have been developed in the context of image-related tasks (Bourtoule et al. 2020; Ginart et al. 2019; Huang et al. 2021; Gao et al. 2023; Nguyen et al. 2022), among others. While machine unlearning is designed to protect the privacy of target samples, (Chen et al. 2021) demonstrates that in the context of model classification tasks, machine unlearning might leave traces.
@@
Reinforcement Learning from Human Feedback: Numerous studies have explored using human feedback to optimize models in various settings, such as simulated robotic control (Christiano et al. 2017), game-playing (Bradley Knox and Stone 2008), machine translation (Nguyen, Daumé III, and Boyd-Graber 2017), citation retrieval (Menick et al. 2022), browsing-based question-answering (Song, Meng, and Ermon 2021), summarization (Stiennon et al. 2020; Ziegler et al. 2020), instruction-following (Ouyang et al. 2022), and alignment with specifications (Bai et al. 2022). Recently, (Lee et al. 2023) studied the alignment of text-to-image diffusion models to human preferences using a method based on reward-weighted likelihood maximization. Their method corresponds to one iteration of the reward-weighted regression (RWR) method. Additionally, (Black et al. 2023a) proposed a class of policy gradient algorithms to perform reinforcement learning by posing denoising diffusion as a multi-step decision-making problem. Their findings show that DDPO significantly outperforms multiple iterations of weighted likelihood maximization (RWR-style) optimization. Therefore, we adopt the DDPO method for reinforcement learning in this context.
@@ 3 Problem Formulation
Considering a training image dataset , composed of both copyrighted () and non-copyrighted () images. The dataset proportions are represented by and , with .
For each image in the dataset, we associate a feature vector and a corresponding text prompt . The diffusion model is trained to generate an image from the prompt , aiming to produce a feature vector that is close to .
To ensure the model does not generate images that infringe on copyrights, we define a copyright loss CL, which penalizes the generation of images that are overly similar to those in . Additionally, we employ the Fréchet Inception Distance (FID) to evaluate the quality of the generated images, ensuring that they are both visually coherent and diverse.
The objective is to train the model in such a way that it minimizes the copyright loss while maintaining a high FID score, effectively balancing the trade-off between reducing copyright infringement and preserving image quality.
@@ 4 Main Approach
4.1 Overview
we propose a novel approach to minimize copyright infringement in text-to-image diffusion models by leveraging reinforcement learning (RL) and our proposed copyright metrics. We first define a copyright metric to measure how closely a generated image resembles copyrighted content. Then, we integrate this copyright metric into reward function and employ reinforcement learning techniques to fine-tune a pre-trained text-to-image diffusion model. Specifically, the model is trained to maximize the reward by iteratively adjusting its parameters to reduce the likelihood of producing copyright-infringing images. By doing so, we ensure that the model maintains high image quality while adhering to copyright constraints.
As shown in Figure 2, the training process of RLCP is as follows:
Gather Datasets: Compile datasets that include both original and copyright-infringing samples.
Prompts Generation: Fed these images into the CLIP interrogator, allowing us to obtain prompts that correspond to each anchor image. The CLIP Interrogator is utilized to convert copyrighted images into corresponding textual information. This text is subsequently refined and transformed into prompts, which are then inputted into a diffusion model to generate the corresponding infringing images.
Model Training: the prompts are used as input for the stable diffusion model, resulting in the generation of images by the stable diffusion model.
Discriminator-Based Scoring: Use a discriminator in the reward model to score generated samples based on the two metrics. Samples less similar to the copyrighted data receive higher rewards.
@@ 4.2 Copyright Metric
Semantic metric. We leverage the CLIP model(Radford et al. 2021) to generate semantic embeddings for anchor and generated images and calculate the metrics by:
@@
Perceptual metric. Here we used the Learned Perceptual Image Patch Similarity (LPIPS) metric proposed by (Zhang et al. 2018). First, LPIPS normalises the feature dimension of all pixels and layers to unit length, scales each feature by a specific weight, and we then calculated squared distance between these weighted activations. The squared distances are then averaged across the image dimensions (spatial averaging) and summed over the layers, resulting in the final perceptual distancemetric as follows:
@@
We normalize the result and subtract it to 1 to obtain a metric for perceptual similarity as follows:
Finally, we denote the total copyright loss as:
@@ 4.3 The proposed method RLCP
RLCP aims to achieve the dual objectives of maintaining high-quality image generation and ensuring compliance with copyright laws, thereby addressing the pressing challenge of copyright infringement in text-to-image diffusion models. It consists of several key components:
Reward-Based Learning Framework: Our approach involves using a discriminator in the reward model to score generated samples. Images that are less similar to copyrighted data receive higher rewards. This reward system helps guide the model towards generating compliant images.
Fine-Tuning with Reinforcement Learning: Starting from a pre-trained text-to-image diffusion model, we fine-tune it using RL techniques to optimize the generation process according to the defined reward functions. This involves framing the denoising process as a multi-step decision-making problem and applying policy gradient algorithms to maximize the reward signal.
KL Regularization: To prevent the model from overfitting to the reward function, we introduce KL divergence as a regularization term. This helps maintain the generative capabilities of the original diffusion model while steering it towards producing non-infringing content.
Specific training process
we describe how we fine-tune a pre-trained text-to-image diffusion model using reinforcement learning (RL) to minimize the risk of generating copyright-infringing content. Our approach is guided by a reward function that leverages the copyright loss (CL) metric defined in Eq. (5), and incorporates KL regularization to ensure the model maintains high-quality image generation.
Model Initialization: We begin with a pre-trained diffusion model , where denotes the model parameters. The model is initialized based on a large-scale text-to-image dataset that includes both copyrighted () and non-copyrighted () data.
Reward Function Design: The reward function is the key component guiding the model’s training. It is defined as a weighted sum of the semantic similarity loss and the perceptual similarity loss :
where:
represents the semantic similarity loss, calculated as the Mean Squared Error (MSE) between the generated image embeddings and the embeddings of the original copyright images.
represents the perceptual similarity loss, measured using the LPIPS metric.
The parameters and are the same as Equation.(5), allowing fine-tuning of the sensitivity to copyright similarity.
Training with DDPO: We utilize Denoising Diffusion Policy Optimization (DDPO) (Black et al. 2023b) to optimize the model. This approach treats the denoising process as a sequence of actions in a Markov Decision Process (MDP). The goal is to maximize the expected reward:
The model iteratively updates the parameters using gradient ascent to increase the reward, which in turn reduces the likelihood of generating infringing content.
KL Regularization: To prevent the model from overfitting to the reward function, we introduce a KL divergence term that regularizes the training. This term penalizes significant deviations from the original model distribution:
The final objective function that the model optimizes combines the reward function with KL regularization:
where is a regularization parameter that balances the trade-off between achieving high rewards (minimizing copyright loss) and maintaining the generative capabilities of the original diffusion model.
@@ 5 Experiment
In this section, we first evaluate the effectiveness of our proposed copyright loss metric. Then we evaluate RLCP on 3 real-world datasets. Furthermore, we explore the impact of the proportion of copyright images in the training set on the efficiency of RLCP.
@@ 5.1 Experiment Setup
Paintings: Painting artworks(WikiArt 2024) often embody the distinctive style of the artist, encompassing aspects such as brushstrokes, lines, colors, and compositions. We gathered over 1000 paintings from Vincent Van Gogh.
@@
Cartoon Figures: Cartoon figure images, including characters from animations and cartoons, are often protected by law. Similar to portraits, we have curated a dataset of around 1000 influential animated characters and figures by collecting information from reputable sources such as Kaggle222https://www.kaggle.com/datasets and Wikipedia333https://www.tensorflow.org/datasets/catalog/wikipedia.
@@
Portrait: The right of a portrait encompasses an individual’s authority over their own image, including their facial features, likeness, and posture. In our research, we gathered over 500 portrait images from Wikipedia, an open-access, multilingual online encyclopedia that provides information on a multitude of subjects.
For the non-copyright dataset, we sourced images from ImageNet (Deng et al. 2009), selecting one image from each class, resulting in a total of 1,000 images.
Evaluation Metric. We use our proposed metric Copyright Loss (CL) as well as CLIP and norm to measure the degree of copyright violation. We also use FID to measure the generative quality of text-to-image diffusion model.
@@
CLIP: We evaluate changes of CLIP scores, for text-image similarity.
@@
Copyright Loss (CL): We quantify the similarity between the original copyright images and their unlearned counterparts on the feature level, utilizing our proposed CL(copyright loss) metric.
@@
Fréchet Inception Distance (FID): We use FID to evaluate the generative image quality of text-to-image diffusion model. The formulation of the metric is as follows:
(6) where:
and represent the mean and covariance matrix of the feature vectors from real images.
and represent the mean and covariance matrix of the feature vectors from generated images.
denotes the trace of a matrix.
@@
Stable diffusion (SDXL): Stable Diffusion(von Platen et al. 2022) is a latent text-to-image diffusion model capable of generating photo-realistic images given any text input.
Forget-Me-Not: Forget-Me-Not(Zhang et al. 2023) is an efficient and low-cost solution designed to safely remove specified IDs, objects, or styles from a well-configured text-to-image model in as little as 30 seconds, without impairing its ability to generate other content.
Concept Removal (CA): CA(Gandikota et al. 2023) An efficient method of ablating concepts in the pretrained model.
Unified Concept Editing(UCE): UCE(Kumari et al. 2023) edits the model without training using a closed-form solution, and scales seamlessly to concurrent edits on text-conditional diffusion models.
Experimental environment and hyperparameters. All the experiments were conducted on a cluster with 2 80Gb A100 GPUs. The hyperparameters are listed in Table 1.
| batch_size | 8 |
| learning_rate | 3 |
| Samples per iteration | 32 |
| Gradient updates per iteration | 4 |
| Clip range |
| Metric | SD | RLCP | CA | UCE | Forget-Me-Not |
| CLIP | 65.78 | 35.12 | 50.34 | 52.11 | 48.76 |
| CL | 55.24 | 28.45 | 45.23 | 48.12 | 40.67 |
| -norm | 0.62 | 0.40 | 0.52 | 0.55 | 0.47 |
| FID | 13.4 | 18.1 | 14.9 | 14.6 | 15.1 |
| Metric | SD | RLCP | CA | UCE | Forget-Me-Not |
| CLIP | 75.56 | 30.34 | 65.45 | 67.89 | 60.33 |
| CL | 70.23 | 20.76 | 60.89 | 63.12 | 55.47 |
| -norm | 0.72 | 0.33 | 0.62 | 0.65 | 0.57 |
| FID | 15.5 | 20.8 | 16.7 | 16.2 | 17.0 |
| Metric | SD | RLCP | CA | UCE | Forget-Me-Not |
| CLIP | 60.12 | 30.23 | 48.34 | 50.89 | 45.56 |
| CL | 52.89 | 25.67 | 42.23 | 44.78 | 39.12 |
| -norm | 0.58 | 0.38 | 0.46 | 0.49 | 0.43 |
| FID | 12.7 | 17.5 | 14.3 | 14.1 | 14.8 |

### v3

@@ Abstract
The increasing sophistication of text-to-image generative models raises challenges in defining and enforcing copyright criteria. Existing methods like watermarking and dataset deduplication fall short due to the lack of standardized metrics and the complexity of addressing copyright issues in diffusion models. To tackle these challenges, we propose RLCP, a Reinforcement Learning-based Copyright Protection method for Text-to-Image Diffusion Models. Our approach introduces a novel copyright metric grounded in legal precedents and employs the Denoising Diffusion Policy Optimization (DDPO) framework to minimize copyright-infringing content while preserving image quality. A reward function based on our metric and KL divergence regularization ensures stable fine-tuning. Experiments on mixed datasets of copyright and non-copyright images show that RLCP effectively reduces copyright infringement risk without compromising output quality.
Index Terms— Copyright Protection, Diffusion Model, Reinforcement Learning
@@ 1 Introduction
Recently, text-to-image diffusion models have garnered significant attention in research. These advanced methods [1, 2] have demonstrated exceptional capabilities in converting textual descriptions into highly accurate and visually coherent images. The advancements in these techniques have unlocked numerous possibilities for various downstream tasks, including image editing [3, 4], image denoising [4, 5], and super-resolution [6].
While the progress in text-to-image generative models has profoundly impacted different industries, it also presents significant challenges for copyright protection. These models utilize extensive training data that may include copyrighted works, which they are sometimes capable of memorizing[7]. This ability can result in the production of images that closely resemble protected content in Figure 1, posing significant challenges to copyright protection[8]. Recent legal cases, such as those involving Stable Diffusion[9] and Midjourney[10], highlight concerns over the use of copyrighted data in AI training, where the models potentially infringe on the rights of numerous artists. These cases highlight a growing concern: Could the high-quality content synthesized by these generative AIs be excessively similar to copyrighted training data, potentially violating the rights of copyright holders?
Various methods have been proposed for source data copyright protection. One approach involves using unrecognizable examples[11, 12] that prevent models from learning key features of protected images either during inference or training stages. However, this method is highly dependent on the specific image and model, and it lacks general reliability. Machine unlearning[13, 14, 15] removes contributions of copyright data, aligning with the right to be forgotten, while dataset deduplication[16] helps reduce the risk of training sample memorization.
@@
To tackle these challenges, we propose a Reinforcement Learning-based Copyright Protection method (RLCP) for text-to-image diffusion models to reduce the possibility of generating copyright-infringing content. Specifically, inspired by Courts111https://law.justia.com/cases/federal/appellate-courts/F2/562/1157/293262/ in the US which employs two-part test to determine copyright violation, which contains an extrinsic test examining objective similarity in specific expressive elements, and an intrinsic test assessing subjective similarity from the perspective of a reasonable audience, we first propose copyright metric that mirror these legal standards by combining semantic and perceptual similarity. Then, we propose a novel framework that combines reinforcement learning with copyright infringement metrics. We leverage the Denoising Diffusion Policy Optimization framework to guide the model through a multi-step decision-making process, optimizing it using a reward function that incorporates our proposed copyright metric. Additionally, we employ KL divergence as a regularization term to mitigate some failure modes and stabilize RL fine-tuning.
@@ 2 Related Work
Text-to-Image Diffusion Models: Recently, text-to-image diffusion models have garnered significant attention in research. These advanced methods [1, 2] have demonstrated exceptional capabilities in converting textual descriptions into visually coherent and realistic images with high accuracy. The advancements in these techniques have unlocked numerous possibilities for various downstream tasks, including image editing [3, 4], image denoising [4, 5], and super-resolution [6].
@@
Copyright Protection: Several studies in the legal literature have examined copyright issues in machine learning and data mining, focusing primarily on potential infringements during the training phase.: (1) Concept Removal: To remove explicit artwork from large models, [11] presents a fine-tuning method for concept removal from diffusion models. Additionally, [12] presents the ”Forget-Me-Not” method, which enables the targeted removal of specific objects and content from large models within 30 seconds while minimizing the impact on other content. (2) Dataset Deduplication: [16] explores whether diffusion models create unique artworks or directly replicate certain content from the training dataset during image generation. (3) Machine Unlearning: Numerous model unlearning methods have been developed in the context of image-related tasks [13, 14, 15], among others. While machine unlearning is designed to protect the privacy of target samples, [17] demonstrates that in the context of model classification tasks, machine unlearning might leave traces.
@@
Reinforcement Learning from Human Feedback: Numerous studies have explored using human feedback to optimize models in various settings, such as simulated robotic control [18], game-playing [19], machine translation [20], citation retrieval [21], browsing-based question-answering [22], summarization [23], instruction-following [24], and alignment with specifications [25]. Recently, [26] studied the alignment of text-to-image diffusion models to human preferences using a method based on reward-weighted likelihood maximization. Their method corresponds to one iteration of the reward-weighted regression (RWR) method. Additionally, [27] proposed a class of policy gradient algorithms to perform reinforcement learning by posing denoising diffusion as a multi-step decision-making problem. Their findings show that DDPO significantly outperforms multiple iterations of weighted likelihood maximization (RWR-style) optimization. Therefore, we adopt the DDPO method for reinforcement learning in this context.
@@ 3 Problem Formulation
We consider a training dataset , composed of copyrighted () and non-copyrighted () images. The dataset proportions are represented by and , with .
Each image in the dataset has an associated feature vector and a text prompt . The model generates an image from , aiming for to closely match .
To prevent copyright violations, we define a copyright loss (CL) that penalizes generated images that are overly similar to those in . Additionally, the Fréchet Inception Distance (FID) evaluates the quality of generated images to ensure they are visually coherent and diverse.
The goal is to train the model to minimize the copyright loss (CL) while maintaining a high FID score, effectively balancing the trade-off between reducing copyright infringement and preserving generative image quality.
@@ 4.1 Copyright Metric
Semantic metric. We leverage the CLIP model[28] to generate semantic embeddings for anchor and generated images and calculate the metrics by:
@@
Perceptual metric. Here we used the Learned Perceptual Image Patch Similarity (LPIPS) metric proposed by [29]. First, LPIPS normalizes the feature dimension of all pixels and layers to unit length, scales each feature by a specific weight, and we then calculated squared distance between these weighted activations. The squared distances are then averaged across the image dimensions (spatial averaging) and summed over the layers, resulting in the final perceptual distance metric as follows:
@@
We normalize the result and subtract it to 1 to obtain a metric for perceptual similarity and semantic similarity as follows:
Finally, we denote the total copyright loss between a query image y and value image x as:
@@ 4.2 The proposed method RLCP
RLCP aims to achieve the dual objectives of maintaining high-quality image generation and ensuring compliance with copyright laws, thereby addressing the pressing challenge of copyright infringement in text-to-image diffusion models. The overview progress shows in 2, it consists of several key steps:
4.2.1 Model Initialization
We start with a pre-trained text-to-image diffusion model , where represents the parameters trained on a dataset containing both copyrighted () and non-copyrighted () data. Diffusion models like DDPM approximate the data distribution through a reverse Markov chain modeled as:
where the forward diffusion process adds Gaussian noise, and the reverse process iteratively denoises to recover .
4.2.2 Reward Function Design
4.2.3 Training with DDPO
The training process leverages Denoising Diffusion Policy Optimization (DDPO), framing the reverse diffusion process as a Markov Decision Process (MDP). The model aims to maximize the expected reward:
Policy gradient methods iteratively adjust to increase , reducing the likelihood of generating infringing content while preserving image quality.
4.2.4 KL Regularization
To prevent overfitting to the reward function, KL divergence regularization is added:
penalizing deviations from the original model distribution. The final loss function combines reward optimization with KL regularization:
where controls the trade-off between compliance and generative fidelity.
@@ 5 Experiment
In this section, we first evaluate the effectiveness of our proposed copyright loss (CL) metric. Then we evaluate RLCP on 3 real-world datasets. Furthermore, we explore the impact of the proportion of copyright images in the training set on the efficiency of RLCP.
@@ 5.1 Experiment Setup
Paintings: Painting artworks[30] often embody the distinctive style of the artist, encompassing aspects such as brushstrokes, lines, colors, and compositions. We gathered over 1000 paintings from Vincent Van Gogh.
@@
Cartoons: Cartoon figure images, including characters from animations and cartoons, are often protected by law. We have curated a dataset of around 1000 influential animated characters and figures by collecting information from reputable sources such as Kaggle222https://www.kaggle.com/datasets and Wikipedia333https://www.tensorflow.org/datasets/catalog/wikipedia.
@@
Portraits: The right of a portrait encompasses an individual’s authority over their own image, including their facial features, likeness, and posture. We gathered over 500 portrait images from Wikipedia.
For the non-copyright dataset, we sourced images from ImageNet [31], selecting one image from each class, resulting in a total of 1,000 images.
| Method | Paintings | Cartoons | Portraits | |||||||||
| CLIP (%) | CL (%) | -norm (%) | FID | CLIP (%) | CL (%) | -norm (%) | FID | CLIP (%) | CL (%) | -norm (%) | FID | |
| SDXL | 65.78 | 55.24 | 62.10 | 13.4 | 75.56 | 70.23 | 72.21 | 15.1 | 60.12 | 52.89 | 58.31 | 12.7 |
| RLCP | 35.12 | 28.45 | 40.12 | 18.1 | 30.34 | 20.76 | 33.21 | 20.8 | 30.23 | 25.67 | 38.34 | 17.5 |
| CA | 50.34 | 45.23 | 52.40 | 14.9 | 65.45 | 60.89 | 62.23 | 16.7 | 48.34 | 42.23 | 46.12 | 14.3 |
| UCE | 52.11 | 48.12 | 56.01 | 14.6 | 67.89 | 63.12 | 65.46 | 16.2 | 50.89 | 44.78 | 49.34 | 14.1 |
| Forget-Me-Not | 48.76 | 40.67 | 47.42 | 15.1 | 60.33 | 55.47 | 57.34 | 17.0 | 45.56 | 39.12 | 43.27 | 14.8 |
Evaluation Metric. We use our proposed metric copyright loss (CL) as well as CLIP and -norm to measure the degree of copyright violation. We also use FID to measure the generative quality of text-to-image diffusion model.
@@
CLIP: We evaluate CLIP score changes for text-image similarity.
@@
CL: We quantify the similarity between the original copyright images and their unlearned counterparts utilizing our proposed copyright loss (CL) metric.
@@
Fréchet Inception Distance (FID): We use FID to evaluate the generative image quality of text-to-image diffusion model.
@@
SDXL: SDXL[32] is a latent text-to-image diffusion model capable of generating photo-realistic images given any text input.
Concept Removal (CA): CA[11] An efficient method of ablating concepts in the pretrained model.
Unified Concept Editing (UCE): UCE[33] edits the model without training using a closed-form solution, and scales seamlessly to concurrent edits on text-conditional diffusion models.
Forget-Me-Not: Forget-Me-Not[12] is an efficient solution designed to safely remove specified IDs, objects, or styles from a well-configured text-to-image model.
Experimental environment and hyperparameters. All the experiments were conducted on a cluster with 2 80Gb A100 GPUs. The hyperparameters used were: a batch size of 8, a learning rate of , 32 samples per iteration, 4 gradient updates per iteration, and a clip range of .